OpenStudy (anonymous):

lim x->0 (1+tan(x)^1/x

OpenStudy (whpalmer4):

I assume that is supposed to be \[\lim_{x \rightarrow 0} (1+\tan(x))^{1/x}\] Often a good thing to explore is changing it into some sort of fraction and using L'Hopital's rule. You could take the natural log and raise e to the resulting log which would give you \[\huge{{\lim_{x\rightarrow 0}(e^{\frac{\ln (1+\tan(x)) }{x}})}}\] and looks promising.

OpenStudy (anonymous):

i'm really bad with logarithms, etc.. but i'm guessing you multiplied by e and ln because they cancel each other out, then by multiplying with ln, the exponent can be "taken down" and multiplied by the function?

OpenStudy (whpalmer4):

No, I was taking advantage of the fact that \[e^{\ln x} = x\]

OpenStudy (anonymous):

\[\begin{align*}\left(1+\tan x\right)^{1/x}&=\exp\bigg[\ln(1+\tan x)^{1/x}\bigg]\\ &=\exp\bigg[\frac{1}{x}\ln(1+\tan x)\bigg]\\ &=\exp\bigg[\frac{\ln(1+\tan x)}{x}\bigg] \end{align*}\] as per @whpalmer4's suggestion. If you're not familiar with the notation, \[\large e^{f(x)}=\exp[f(x)]\] Then, using the continuity of the exponential function at 0, you can change the limiting function: \[\lim_{x\to0}\exp\bigg[\frac{\ln(1+\tan x)}{x}\bigg]=\exp\bigg[\lim_{x\to0}\frac{\ln(1+\tan x)}{x}\bigg]\]

OpenStudy (whpalmer4):

Yeah, that's much easier to read with the exp notation, thanks!

OpenStudy (whpalmer4):

after you apply L'Hopital, the fraction ought to reduce to something you can directly evaluate at \(x=0\) and then you've just got \(e\) to that power as your limit.

OpenStudy (whpalmer4):

Oh, also \[\ln u^n = n \ln u\]makes an appearance to generate that fraction: \[u = 1+\tan x\]and\[n =1/x\]

OpenStudy (anonymous):

thanks! But how come i can change the limit function?

OpenStudy (anonymous):

and after i get limx→0exp[ln(1+tanx)x] i can apply lhopital?

OpenStudy (jhannybean):

I don't understand the "exp" notation... What is exp?

OpenStudy (whpalmer4):

@Jhannybean \[\exp(x) = e^{x}\]

OpenStudy (anonymous):

it's just different notation

OpenStudy (jhannybean):

Oh okay.

OpenStudy (whpalmer4):

Thanks to my web browser for discarding a nearly complete post without confirmation :-( We had as the guts of our limit \[(1+\tan x)^{1/x}\] which we can rewrite without changing its value as\[(1/x) \ln (1+\tan x)\] or\[\frac{\ln(1+\tan x)}{x}\] The reason we can do that is thanks to two properties of logarithms: \[\exp[{\ln u}] = u\]\[\ln u^n = n \ln u\] Finally, the limit of a constant (e) to a power is equal to the constant to the limit of the power, so our limit becomes \[\exp[\lim_{x->0} \frac{1+\tan x}{x} ]\]and we can use L'Hopital on the fraction.

OpenStudy (whpalmer4):

\[\frac{d}{dx}[1+\tan x] = \sec^2 x\]\[\frac{d}{dx}[x] = 1\] so our limit after L'Hoping is \[\exp[\lim_{x->0} \sec^2x]\]which can be evaluated directly.

OpenStudy (anonymous):

alright, thank you all for your help

OpenStudy (whpalmer4):

What do you get for your answer?

OpenStudy (anonymous):

hmm, where did ln go?

OpenStudy (anonymous):

exp[limx−>0 1+tanx/x], after exp[ln 1+tanx/x] ?

OpenStudy (anonymous):

because of exp[lnu]=u this propery?

OpenStudy (anonymous):

property*

OpenStudy (anonymous):

but then the exp would also go away..

OpenStudy (whpalmer4):

Oh, sorry, I oversimplified there, didn't I :-) \[\frac{d}{dx}[\ln (1+\tan x)] = \frac{\sec^2x}{1+\tan x}\]by the chain rule.

OpenStudy (anonymous):

oh right, forgot about chain rule

OpenStudy (anonymous):

i'm still confused, which step is that?

OpenStudy (whpalmer4):

doing L'Hopital....

OpenStudy (whpalmer4):

that's the derivative of the numerator, and the derivative of the denominator is just 1

OpenStudy (whpalmer4):

in any case, sec^2 x/(1+tan x) can be evaluated at x = 0

OpenStudy (anonymous):

i understand everything to: exp {ln(1+tanx)/x} Using: exp {ln u} = u I should get: (1+tanx)/x , right?

OpenStudy (anonymous):

oh wait, u= 1+tanx, i get it

OpenStudy (anonymous):

but i would think my 'u' changed to 1+tanx/x ....

OpenStudy (whpalmer4):

remember, the limit is inside the exp[ ]...

OpenStudy (jhannybean):

\(\large \cfrac{\ln(1+\tan(x))}{x}\) using LH rule : \(\large \cfrac{f'(x)}{g'(x)}\)\[\large \cfrac{\cfrac{1\cdot \sec^2(x)}{1+\tan(x)}}{1} = \cfrac{\sec^2(x)}{1+\tan(x)}\]

OpenStudy (whpalmer4):

so we take the derivative of ln (1+tan x) over the derivative of x as we do L'Hopital's rule, and that gives us sec^2 x/(1+tan x) which we evaluate at x = 0. tan (0) = 0, sec (0) = 1, so we have exp[1^2/(1+0)] = exp[1] = ...

OpenStudy (jhannybean):

And @whpalmer4 just explained word per word my evaluation of the function. :)

OpenStudy (anonymous):

sp the answer should just be e

OpenStudy (anonymous):

so*

OpenStudy (whpalmer4):

yep, eeeeeeeee the magical number.....

OpenStudy (anonymous):

and i use the quotient rule to derive

OpenStudy (whpalmer4):

quotient rule? chain rule, isn't it? you're taking the derivative of ln u where u = 1 + tan x

OpenStudy (anonymous):

i see

OpenStudy (jhannybean):

\[\huge e^{\lim_{x \rightarrow 0}\cfrac{\sec^2(x)}{1+\tan(x)}} = e^{\cfrac{\sec^2(0)}{1+\tan(0)}} = e^{\cfrac{1}{1}}=e\]

OpenStudy (jhannybean):

lol. hideous.....but readable.

OpenStudy (anonymous):

nice, thanks